A company borrows $25,000 from a bank, with an interest rate of 3%. The terms of borrowing provide that interest accrues based on annual compounding, and the principle and interest are due in full 4 years from the date of borrowing. What is the total amount due on this borrowing at the due date?
Answer #1

Single Payment Compound Amount Factor (SPCAF)=(F/P,i,N)=(1+i)^N

i=Annual compounding interest rate =3%=0.03

N=Number of Years=4

SPCAF=(F/P,3%,4)=(1+0.03)^4=1.125509

Total Amount Due on the borrowing at end of 4 years=$25,000*1.125509=$28,137.72

Total amount due $28,137.72